# Existence of attracting periodic orbits in 3-dimensional strongly 2-cooperative systems

### v1

@@ II Notation and Preliminary Results
The dynamical system (3), having the set of equilibria $\mathcal{E}$, satisfies the strong Poincaré-Bendixson property if, for any bounded solution $x(t,a)$, with $a\in\Omega$, it holds that
$\omega(a)\cap\mathcal{E}=\emptyset\implies\omega(a)\ \text{is a periodic orbit}.$
Intuitively, this suggests that all solutions behave like the solutions of a planar dynamical system.
Since a strongly $2$-cooperative system satisfies the strong Poincaré-Bendixson property [35], establishing strong $2$-cooperativity provides important information on the asymptotic behaviour of the non-linear system.
The analysis of sign patterns guaranteeing $k$-positivity and $k$-cooperativity relies on the theory of compound matrices (see, e.g., [4]), a fundamental tool also for $k$-contractive [37], $\alpha$-contractive [38], and totally positive differential systems [20]. A sign pattern related to (2) appears in the seminal work by Mallet-Paret and Smith on monotone cyclic feedback systems [19] (see also [5, 10, 34] and the references therein for more recent results).
Instability of a $3\times 3$ strongly $2$-cooperative matrix. A matrix $A\in{\mathbb{R}}^{n\times n}$ with a strongly $k$-positive sign pattern has special spectral properties [1]. Here, we state these properties in the special case with $n=3$, $k=2$, and $A$ is unstable, that is, it admits an eigenvalue with a positive real part.
Lemma 1.
Given $A\in{\mathbb{R}}^{3\times 3}$, consider its characteristic polynomial
$P_{A}(s):=\det(sI_{3}-A)=s^{3}+a_{2}s^{2}+a_{1}s+a_{0}.$
If the free coefficient $a_{0}$ is positive and $A$ is unstable, then $A$ admits one negative real eigenvalue, and two eigenvalues with a positive real part. If, in addition, the system $\dot{x}=Ax$ is strongly $2$-positive, then the eigenvector $\zeta\in{\mathbb{R}}^{3}$ corresponding to the negative real eigenvalue has sign pattern either $\begin{bmatrix}+&-&+\end{bmatrix}^{\top}$ or $\begin{bmatrix}-&+&-\end{bmatrix}^{\top}$, so
$s^{-}(\zeta)=2.$ (5)
Consider the system $\dot{x}=Ax$ and let $x(t)$ be a solution. Lemma 1 implies that $x(t)$ converges to the origin iff $x(t)\in\operatorname{span}(\zeta)$ for any $t$, meaning that $\zeta$ is the only “direction of convergence” to the equilibrium for the system.
Proof:
Let $\lambda_{i}$, $i=1,2,3$, denote the eigenvalues of $A$. Since $P_{A}(s)$ has real coefficients and is of odd degree, $A$ has at least one real eigenvalue. Moreover, since $A$ is unstable, one of its eigenvalues has a positive real part. Since $\lambda_{1}\lambda_{2}\lambda_{3}=-a_{0}<0$, matrix $A$ must admit one negative real eigenvalue, and the other two eigenvalues must be either both positive real, or a complex-conjugate pair with a positive real part.
The determinant of a $2\times 2$ submatrix of a matrix $B\in{\mathbb{R}}^{n\times n}$ is a $2\times 2$ minor of $B$. If $\dot{x}=Ax$ is strongly 2-positive, then all the $2\times 2$ minors of $\exp(A)$ are positive [35]. The eigenvalues of $\exp(A)$ are $\exp(\lambda_{i})$, $i=1,2,3$. Order the $\lambda_{i}$s so that $|\exp(\lambda_{1})|\geq|\exp(\lambda_{2})|\geq|\exp(\lambda_{3})|$. Then, [1, Thm. 2] implies that the product $\exp(\lambda_{1})\exp(\lambda_{2})$ is real and positive, and that the eigenvalues of $\exp(A)$ satisfy the spectral gap condition
$|\exp(\lambda_{2})|>|\exp(\lambda_{3})|.$ (6)
Hence, $\lambda_{3}$ is the real and negative eigenvalue. Using [1, Thm. 2] again gives that the eigenvector $\zeta$ of matrix $\exp(A)$ corresponding to the eigenvalue $\exp(\lambda_{3})$ satisfies (5), which completes the proof. ∎
The next simple example demonstrates Lemma 1.
Example 1.
Let $A:=\begin{bmatrix}-1&0&-9\\ 1&-1&0\\ 0&1&-1\end{bmatrix}$. The system $\dot{x}=Ax$ is strongly $2$-positive, because $A$ is irreducible and has the sign pattern $\bar{A}_{2}$. We can compute numerically $\exp(A)=\begin{bmatrix}-0.1433&-1.4137&-2.1219\\ 0.2358&-0.1433&-1.4137\\ 0.1571&0.2358&-0.1433\end{bmatrix}$ (all numerical values in this paper are to four-digit accuracy), as well as its $2\times 2$ minors: ($0.3538$, $0.7028$, $1.6946$, $0.1883$, $0.3538$, $0.7028$, $0.0781$, $0.1883$, $0.3538$), which are all positive.
The characteristic polynomial of matrix $A$ is $P_{A}(s)=\det(sI_{3}-A)=s^{3}+3s^{2}+3s+10$, with a positive free coefficient. Its eigenvalues are $\lambda_{1}=\mu+\jmath\eta$,  $\lambda_{2}=\mu-\jmath\eta$, and $\lambda_{3}=-3.0801$, where $\mu=0.0400$, $\eta=1.8014$ and $\jmath$ is the imaginary unit. Hence, $A$ is unstable. The eigenvector corresponding to the negative real eigenvalue $\lambda_{3}$ is $\zeta=\begin{bmatrix}-0.8823&0.4242&-0.2039\end{bmatrix}^{\top}$, with $s^{-}(\zeta)=2$.
III Existence of an Attracting Periodic Orbit
We can now state and prove our main result on the existence of an attracting periodic orbit in strongly $2$-cooperative 3D systems.
@@ III Existence of an Attracting Periodic Orbit
Suppose that $f\in C^{1}$, and let $J(x):=\frac{\partial}{\partial x}f(x)$ denote the Jacobian of the vector field. Suppose that the system is strongly 2-cooperative, and that its trajectories evolve in the closed box
=${\mathcal{B}}=\{x\in{\mathbb{R}}^{3}\ |\ \underline{x}\leq x\leq\overline{x}\},$
for some $\underline{x},\overline{x}\in{\mathbb{R}}^{3}$ with $\underline{x}\leq\overline{x}$. Suppose also that ${\mathcal{B}}$ contains a unique equilibrium $e\in\operatorname{int}({\mathcal{B}})$ that is unstable, and that the free coefficient in the characteristic polynomial $P(s)=\det(sI_{3}-J(e))$ is positive. Partition ${\mathcal{B}}$ into eight closed sub-boxes:
@@
Then ${\mathcal{B}}_{16}:={\mathcal{B}}_{1}\cup\dots\cup{\mathcal{B}}_{6}$ is an invariant set, and for any initial condition $x(0)\in({\mathcal{B}}_{16}\setminus\{e\})$, the solution of (7) converges to a (non-trivial) periodic orbit.
This result implies of course that ${\mathcal{B}}_{16}$ includes at least one attracting periodic orbit. Since cooperative (that is, 1-cooperative) systems do not admit attracting periodic orbits, it is clear that the result of Thm. 1 cannot be derived using the theory of cooperative systems.
Proof:
Define $z_{i}(t):=x_{i}(t)-e_{i}$, $i=1,2,3$. Then,
@@
For subsequent arguments, it is useful to list the number of sign variations of vectors in these sub-boxes:
$z\in\tilde{\mathcal{B}}_{1}$implies $s^{-}(z)=0$,
$z\in\tilde{\mathcal{B}}_{2}$implies $s^{-}(z)\leq 1$,
$z\in\tilde{\mathcal{B}}_{3}$implies $s^{-}(z)\leq 1$,
$z\in\tilde{\mathcal{B}}_{4}$implies $s^{-}(z)=0$,
$z\in\tilde{\mathcal{B}}_{5}$implies $s^{-}(z)\leq 1$,
$z\in\tilde{\mathcal{B}}_{6}$implies $s^{-}(z)\leq 1$,
$z\in\tilde{\mathcal{B}}_{7}$implies $s^{+}(z)=2$,
$z\in\tilde{\mathcal{B}}_{8}$implies $s^{+}(z)=2$.
Since the system is strongly 2-cooperative, Remark 2 implies that
@@
Hence, ~$\tilde{\mathcal{B}}_{16}:=\tilde{\mathcal{B}}_{1}\cup\dots\cup\tilde{\mathcal{B}}_{6}$ is an invariant set for (8). Note that the equilibrium 00 is contained in ~$\tilde{\mathcal{B}}_{16}$. Hence, a priori, a solution evolving in ~$\tilde{{\mathcal{B}}}_{16}$ may still converge to the equilibrium. Our goal now is to build an invariant set for the dynamics (8) by cutting out from ~$\tilde{\mathcal{B}}_{16}$ a cylinder that contains the origin.
To do this, we first analyze the behaviour near the equilibrium. Let $A$ denote the Jacobian of (8) at the origin. Since the origin is unstable, $A$ is unstable, and Lemma 1 implies that $A$ has two eigenvalues $\lambda_{1},\lambda_{2}$ with a positive real part, and one real eigenvalue $\lambda_{3}<0$. Furthermore, the shifted eigenvector
$\tilde{\zeta}:=\zeta-e$
corresponding to $\lambda_{3}$ satisfies
$(\operatorname{span}(\tilde{\zeta})\setminus\{0\})\cap\tilde{{\mathcal{B}}}_{16}\subseteq\operatorname{int}(\tilde{\mathcal{B}}_{7}\cup\tilde{\mathcal{B}}_{8}).$
In particular, each nonzero solution $z(t)$ of (8) that converges to the origin must satisfy $\lim_{t\to\infty}\operatorname{dist}\left(z(t),\operatorname{span}(\tilde{\zeta})\right)=0$.
In view of Lemma 1, there exists a non-singular $T\in\mathbb{R}^{3\times 3}$ such that
$T^{-1}AT=\small\begin{bmatrix}\lambda_{3}&0&0\\ 0&u_{1}&-v\\ 0&v&u_{2}\end{bmatrix},$ (10)
with $\lambda_{3}<0$, $u_{1},u_{2}>0$, where either (1) $v=0$, so $u_{1},u_{2}$ are two positive real eigenvalues; or (2) $u_{1}=u_{2}$, so $u_{1}\pm jv$ are a complex conjugate pair of unstable eigenvalues.
Let $\angle(a,b)$ denote the angle between the vectors $a,b\in{\mathbb{R}}^{3}$. Since $\tilde{\zeta}\in\operatorname{int}\left(\tilde{{\mathcal{B}}}_{7}\cup\tilde{{\mathcal{B}}}_{8}\right)$, there exists $\xi\in(0,1)$ such that, for all $z\in{\mathcal{B}}\setminus\{0\}$, we have
$|\cos$ $(\angle(T^{-1}z,T^{-1}\tilde{\zeta}))|>1-\xi$
$\implies T^{-1}z\in\operatorname{int}(T^{-1}(\tilde{{\mathcal{B}}}_{7}\cup\tilde{{\mathcal{B}}}_{8})).$
By (10), $T^{-1}\tilde{\zeta}=\begin{bmatrix}1&0&0\end{bmatrix}^{\top}$, so letting $q:=T^{-1}z$ yields
$|\cos(\angle(q,\begin{bmatrix}1&0&0\end{bmatrix}^{\top}))|>1-\xi\implies q\in\operatorname{int}(T^{-1}(\tilde{{\mathcal{B}}}_{7}\cup\tilde{{\mathcal{B}}}_{8})),$ (11)
@@
$h(q)=\small\begin{bmatrix}\lambda_{3}&0&0\\ 0&u_{1}&-v\\ 0&v&u_{2}\end{bmatrix}q+\begin{bmatrix}g_{1}(q)\\ g_{2}(q)\\ g_{3}(q)\end{bmatrix},$ (13)
@@
where the functions $g_{i}(q)$, $i=1,2,3$, are smooth on the compact set $T^{-1}\tilde{\mathcal{B}}$, and there exists $M>0$ such that
$\max_{q\in T^{-1}\tilde{\mathcal{B}}}\max_{i}\frac{\left|g_{i}(q)\right|}{|q|^{2}}\leq M.$ (14)
Define $V:{\mathbb{R}}^{2}\to{\mathbb{R}}_{\geq 0}$ by
$V(p_{1},p_{2}):=(u_{1}^{-1}p_{1}^{2}+u_{2}^{-1}p_{2}^{2})/2.$
@@
We claim that there exists $\eta_{*}>0$ such that for all $0<\eta<\eta_{*}$ the closed set
~$\tilde{\mathcal{H}}_{\eta}:=\{q\in T^{-1}\tilde{\mathcal{B}}_{16}\,|\,V(q_{2},q_{3})\geq\eta\},$ (15)
is an invariant set of (12). To prove this, Let $\eta>0$ and fix an initial condition $q(0)\in\tilde{\mathcal{H}}_{\eta}$. Note that (12) implies that
$\dot{V}(q_{2}(t),q_{3}(t))$ $=u_{1}^{-1}q_{2}(t)\dot{q}_{2}(t)+u_{2}^{-1}q_{3}(t)\dot{q}_{3}(t)$
$=q_{2}^{2}(t)+q_{3}^{2}(t)+s(t),$ (16)
where $s(t):=(u_{2}^{-1}-u_{1}^{-1})vq_{2}(t)q_{3}(t)+u_{1}^{-1}q_{2}(t)g_{2}(q(t))+u_{2}^{-1}q_{3}(t)g_{3}(q(t))$. Since either $v=0$ (in the case of two real positive eigenvalues), or $u_{1}=u_{2}$ (in the case of a complex conjugate pair of unstable eigenvalues), we have
$s(t)=u_{1}^{-1}q_{2}(t)g_{2}(q(t))+u_{2}^{-1}q_{3}(t)g_{3}(q(t)).$
Since $q(t)\in T^{-1}\tilde{{\mathcal{B}}}_{16}$, Eq. (11) implies that
$|\cos(\angle(q(t),\begin{bmatrix}1&0&0\end{bmatrix}^{\top}))|\leq 1-\xi,$ (17)
@@
that is, $\frac{q_{1}^{2}(t)}{|q(t)|^{2}}\leq(1-\xi)^{2}$, and since $\xi\in(0,1)$, this gives $|q(t)|^{2}\leq\frac{1}{1-(1-\xi)^{2}}(q_{2}^{2}(t)+q_{3}^{2}(t))$. Combining this with (14) yields
$|q_{i}(t)g_{i}(q(t))|$ $\leq|q_{i}(t)|\frac{|g_{i}(q(t))|}{|q(t)|^{2}}|q(t)|^{2}$
$\leq|q_{i}(t)|\frac{M}{1-(1-\xi)^{2}}(q_{2}^{2}(t)+q_{3}^{2}(t)),$ (18)
for $i\in\{2,3\}$. Thus,
$|s(t)|$ $\leq\frac{M}{1-(1-\xi)^{2}}\big{(}u_{1}^{-1}(q_{2}^{2}(t)|q_{2}(t)|+q_{3}^{2}(t)|q_{2}(t)|)$
$+u_{2}^{-1}(q_{2}^{2}(t)|q_{3}(t)|+q_{3}^{2}(t)|q_{3}(t)|)\big{)}.$
Combining this with (III) implies that there exists a sufficiently small $\eta_{*}>0$ such that, for all $0<\eta<\eta_{*}$ and $q(t)\in\tilde{\mathcal{H}}_{\eta}$, we have that $\dot{V}(q_{2}(t),q_{3}(t))\geq 0$, and thus ~$\tilde{\mathcal{H}}_{\eta}$ is indeed an invariant set of the dynamics. By the definition of ~$\tilde{\mathcal{H}}_{\eta}$, the equilibrium point 00 is not in ~$\tilde{\mathcal{H}}_{\eta}$ (see Fig. 1).
@@
Summarizing, for some $\eta_{*}>0$ and any $0<\eta<\eta_{*}$, ~$\tilde{\mathcal{H}}_{\eta}$ is a closed invariant set, and for all $q(0)\in\tilde{\mathcal{H}}_{\eta}$ the solution $q(t)$ is bounded and keeps a positive distance from the equilibrium 00 (and thus from any equilibrium). Combining this with the strong Poincaré-Bendixson property of the system implies that any trajectory emanating from ~$\tilde{\mathcal{H}}_{\eta}$ converges to a periodic orbit. Since any $q(0)\in\tilde{{\mathcal{B}}}_{16}\setminus\{0\}$ belongs to $\cup_{0<\eta<\eta_{*}}\tilde{\mathcal{H}}_{\eta}$, this completes the proof of Thm. 1. ∎
IV Case Studies

### v2

@@ 2 Notation and Preliminary Results
System (3), having equilibrium set $\mathcal{E}$, satisfies the strong Poincaré-Bendixson property if, for any bounded solution $x(t,a)$, with $a\in\Omega$, the condition $\omega(a)\cap\mathcal{E}=\emptyset$ implies that $\omega(a)$ is a periodic orbit.
Intuitively, this suggests that all solutions behave like the solutions of a planar dynamical system. Since a strongly $2$-cooperative system satisfies the strong Poincaré-Bendixson property (Weiss and Margaliot, 2021a), establishing strong $2$-cooperativity provides important information on the asymptotic behaviour of the non-linear system.
3 Convergence to a Periodic Orbit
We can now state our main result.
@@ 3 Convergence to a Periodic Orbit
Suppose that $f\in C^{2}$, and let $J(x):=\frac{\partial}{\partial x}f(x)$ denote the Jacobian of the vector field. Suppose that (5) is strongly 2-cooperative, and that its trajectories evolve in the closed box =${\mathcal{B}}=\{x\in{\mathbb{R}}^{3}\ |\ \underline{x}\leq x\leq\overline{x}\},$ for some $\underline{x},\overline{x}\in{\mathbb{R}}^{3}$ with $\underline{x}\leq\overline{x}$. Suppose also that ${\mathcal{B}}$ contains a unique equilibrium $e\in\operatorname{int}({\mathcal{B}})$ that is unstable, and that $\det(J(e))<0$. Partition ${\mathcal{B}}$ into eight closed sub-boxes:
@@
Then ${\mathcal{B}}_{16}:={\mathcal{B}}_{1}\cup\dots\cup{\mathcal{B}}_{6}$ is an invariant set for (5), and for any initial condition $x(0)\in({\mathcal{B}}_{16}\setminus\{e\})$, the corresponding solution of (5) converges to a (non-trivial) periodic orbit.
Since ${\mathcal{B}}_{16}$ has positive measure, the result of Theorem 1 cannot be derived using the theory of cooperative systems or differentially positive systems (Forni and Sepulchre, 2016), as in such systems almost all bounded solutions converge to an equilibrium.
Proving Theorem 1 requires the following lemma.
Lemma 1.
Suppose that $A\in{\mathbb{R}}^{3\times 3}$ is unstable and that $\det(A)<0$. Then $A$ admits one negative real eigenvalue, and two eigenvalues with a positive real part. If, in addition, the system $\dot{x}=Ax$ is strongly $2$-positive, then the eigenvector $\zeta\in{\mathbb{R}}^{3}$ corresponding to the negative real eigenvalue has sign pattern $[+~{}-~{}+]^{\top}$ or $[-~{}+~{}-]^{\top}$, so
$s^{-}(\zeta)=2.$ (6)
Consider the system $\dot{x}=Ax$ and let $x(t)$ be a solution. Lemma 1 implies that $x(t)$ converges to the origin iff $x(t)\in\operatorname{span}(\zeta)$ for all $t$, so the direction of the eigenvector $\zeta$ is the only “direction of convergence” to the equilibrium for the system. Hence, the equilibrium is a saddle point.
Proof of Lemma 1. Since $n=3$, matrix $A$ has at least one real eigenvalue. Also, being $A$ unstable, at least one of its eigenvalues has a positive real part. Let $\lambda_{1},\lambda_{2},\lambda_{3}$ denote the eigenvalues of $A$. Since $\lambda_{1}\lambda_{2}\lambda_{3}=\det(A)<0$, $A$ must admit one negative real eigenvalue, and the other two eigenvalues must be either both positive real, or a complex-conjugate pair with a positive real part.
If $\dot{x}=Ax$ is strongly 2-positive, then all the $2\times 2$ minors of $\exp(A)$ are positive (Weiss and Margaliot, 2021a). The eigenvalues of $\exp(A)$ are $\exp(\lambda_{i})$, $i=1,2,3$. Order the eigenvalues so that $|\exp(\lambda_{1})|\geq|\exp(\lambda_{2})|\geq|\exp(\lambda_{3})|$. Then, (Alseidi et al., 2019, Theorem 2) implies that the product $\exp(\lambda_{1})\exp(\lambda_{2})$ is real and positive, and that the eigenvalues of $\exp(A)$ satisfy the spectral gap condition: $|\exp(\lambda_{2})|>|\exp(\lambda_{3})|.$ Hence, $\lambda_{3}$ is the real and negative eigenvalue. Using (Alseidi et al., 2019, Theorem 2) again gives that the eigenvector $\zeta$ of $\exp(A)$ corresponding to the eigenvalue $\exp(\lambda_{3})$ satisfies (6), which completes the proof.  ∎
Proof of Theorem 1. Define $z_{i}(t):=x_{i}(t)-e_{i}$, $i=1,2,3$. Then,
@@
Note that:
$z\in\tilde{\mathcal{B}}_{1}\cup\tilde{\mathcal{B}}_{4}$implies $s^{-}(z)=0$,
$z\in\tilde{\mathcal{B}}_{2}\cup\tilde{\mathcal{B}}_{3}\cup\tilde{\mathcal{B}}_{5}\cup\tilde{\mathcal{B}}_{6}$implies $s^{-}(z)\leq 1$,
$z\in\tilde{\mathcal{B}}_{7}\cup\tilde{\mathcal{B}}_{8}$implies $s^{+}(z)=2$.
Since the system is strongly 2-cooperative, it follows from Definitions 3 and 4 that
@@
Hence, ~$\tilde{\mathcal{B}}_{16}:=\tilde{\mathcal{B}}_{1}\cup\dots\cup\tilde{\mathcal{B}}_{6}$ is an invariant set for (7). The equilibrium $0\in\tilde{\mathcal{B}}_{16}$, so, a priori, a solution evolving in ~$\tilde{{\mathcal{B}}}_{16}$ may still converge to the equilibrium. Our goal now is to build an invariant set for the dynamics (7) by “cutting out” from ~$\tilde{\mathcal{B}}_{16}$ a cylinder that contains the origin.
To do this, we first analyze the behaviour near the equilibrium. Let $A$ denote the Jacobian of (7) at the origin. Since the origin is unstable, $A$ is unstable, and Lemma 1 implies that $A$ has two eigenvalues $\lambda_{1},\lambda_{2}$ with a positive real part, and one real eigenvalue $\lambda_{3}<0$. Furthermore, the shifted eigenvector $\tilde{\zeta}:=\zeta-e$ corresponding to $\lambda_{3}$ satisfies: $(\operatorname{span}(\tilde{\zeta})\setminus\{0\})\cap\tilde{{\mathcal{B}}}\subseteq\operatorname{int}(\tilde{\mathcal{B}}_{7}\cup\tilde{\mathcal{B}}_{8}).$
As a consequence of Lemma 1, there exists a non-singular matrix $T\in\mathbb{R}^{3\times 3}$ such that, given $\lambda_{3}<0$ and $u_{1},u_{2}>0$,
$T^{-1}AT=\begin{bmatrix}\lambda_{3}&0&0\\ 0&u_{1}&v_{1}/\delta\\ 0&v_{2}\delta&u_{2}\end{bmatrix},$ (9)
where $\delta$ is a scaling parameter that will be chosen subsequently, and three cases are possible: (i) $v_{1}=v_{2}=0$, so $u_{1}=\lambda_{1}$ and $u_{2}=\lambda_{2}$ are two positive real eigenvalues; (ii) $u_{1}=u_{2}$ and $v_{1}/\delta=-v_{2}\delta$, so $\lambda_{1,2}=u_{1}\pm jv_{1}$ are a complex conjugate pair of unstable eigenvalues; (iii) $\lambda_{1}=u_{1}=u_{2}=\lambda_{2}$, $v_{1}=1$ and $v_{2}=0$.
Let $\angle(a,b)$ denote the angle between the vectors $a,b\in{\mathbb{R}}^{3}$. Since, up to scaling, $\tilde{\zeta}\in\operatorname{int}\left(\tilde{{\mathcal{B}}}_{7}\cup\tilde{{\mathcal{B}}}_{8}\right)$, for any fixed $\delta$ there exists $\xi\in(0,1)$ such that, for all $z\in\tilde{\mathcal{B}}\setminus\{0\}$, we have that $|\cos(\angle(T^{-1}z,T^{-1}\tilde{\zeta}))|>1-\xi\implies T^{-1}z\in\operatorname{int}(T^{-1}(\tilde{{\mathcal{B}}}_{7}\cup\tilde{{\mathcal{B}}}_{8}))$.
By (9), $T^{-1}\tilde{\zeta}=[1~{}0~{}0]^{\top}$, so letting $q:=T^{-1}z$ yields
$|\cos(\angle(q,[1~{}0~{}0]^{\top}))|>1-\xi\implies q\in\operatorname{int}(T^{-1}(\tilde{{\mathcal{B}}}_{7}\cup\tilde{{\mathcal{B}}}_{8})),$ (10)
@@
Let us fix $\delta=1$ in cases (i) and (ii) above, while $\delta>\frac{1}{2u}$ in case (iii).
@@
$h(q)=\begin{bmatrix}\lambda_{3}&0&0\\ 0&u_{1}&v_{1}/\delta\\ 0&v_{2}\delta&u_{2}\end{bmatrix}q+\begin{bmatrix}g_{1}(q)\\ g_{2}(q)\\ g_{3}(q)\end{bmatrix},$ (12)
@@
where the non-linear functions $g_{i}(\cdot)$, $i=1,2,3$, are $C^{2}$ on the compact set $T^{-1}\tilde{\mathcal{B}}$, $g_{i}(0)=0$, $\nabla g_{i}(0)=0$, and there exists $M>0$ such that
$\max_{q\in T^{-1}\tilde{\mathcal{B}}}\,\,\max_{i=1,2,3}\,\frac{\left|g_{i}(q)\right|}{|q|^{2}}\leq M.$ (13)
@@
Define $V\colon{\mathbb{R}}^{2}\to{\mathbb{R}}_{\geq 0}$ as $V(p_{1},p_{2}):=(p_{1}^{2}+p_{2}^{2})/2.$ We claim that there exists $\eta_{*}>0$ such that for all $0<\eta<\eta_{*}$ the closed set
~$\tilde{\mathcal{H}}_{\eta}:=\{q\in T^{-1}\tilde{\mathcal{B}}_{16}\,|\,V(q_{2},q_{3})>\eta\}$ (14)
is an invariant set for (11). Since $T^{-1}\tilde{\mathcal{B}}_{16}$ is an invariant set due to (3), it is enough to show that, for any solution $q(t)$ that emanates from $T^{-1}\tilde{\mathcal{B}}_{16}\setminus\left\{0\right\}$ and satisfies $V(q_{2}(0),q_{3}(0))>\eta$, there cannot exist a time $t^{*}>0$ such that $V(q_{2}(t^{*}),q_{3}(t^{*}))\leq\eta$, provided that $\eta>0$ is small enough.
To prove this, we let $\eta>0$ and analyze $\dot{V}(q_{2},q_{3})$ when $q\in T^{-1}\tilde{\mathcal{B}}_{16}\setminus\tilde{\mathcal{H}}_{\eta}$. By (11),
$\dot{V}(q_{2},q_{3})$ $=q_{2}\dot{q}_{2}+q_{3}\dot{q}_{3}$
$=\begin{bmatrix}q_{2}\\ q_{3}\end{bmatrix}^{\top}\begin{bmatrix}u_{1}&\frac{v_{1}}{2\delta}+\frac{\delta}{2}v_{2}\\ \frac{v_{1}}{2\delta}+\frac{\delta}{2}v&u_{2}\end{bmatrix}\begin{bmatrix}q_{2}\\ q_{3}\end{bmatrix}+s,$ (15)
where $s:=q_{2}g_{2}(q)+q_{3}g_{3}(q)$. Our choices of $\delta$ guarantee that the quadratic form in (3) is positive definite, and hence larger than $\kappa V(q_{2},q_{3})$ for some $\kappa>0$.
We now show that $\kappa V(q_{2},q_{3})$ dominates $s$ when $V(q_{2},q_{3})$ is sufficiently small. Pick $q\in T^{-1}\tilde{{\mathcal{B}}}_{16}\setminus\left\{0\right\}$. Eq. (10) gives
$|\cos(\angle(q,[1~{}0~{}0]^{\top}))|\leq 1-\xi,$ (16)
whence $q_{1}^{2}\leq(1-\xi)^{2}|q|^{2}$, or equivalently
$\left(1-\left(1-\xi\right)^{2}\right)q_{1}^{2}\leq(1-\xi)^{2}(q_{2}^{2}+q_{3}^{2})=2(1-\xi)^{2}V(q_{2},q_{3}).$ (17)
@@
Dividing by $(1-\left(1-\xi\right)^{2})$, adding $(q_{2}^{2}+q_{3}^{2})$ on both sides, and using the fact that $\xi\in(0,1)$, gives $|q|^{2}\leq\frac{1}{1-(1-\xi)^{2}}2V(q_{2},q_{3})$. Combining this with (13) yields that for $i\in\{2,3\}$:
$|q_{i}g_{i}(q)|\leq|q_{i}|\frac{|g_{i}(q)|}{|q|^{2}}|q|^{2}\leq|q_{i}|\frac{2MV(q_{2},q_{3})}{1-(1-\xi)^{2}}.$ (18)
Thus,
$|s|\leq\frac{2MV(q_{2},q_{3})}{1-(1-\xi)^{2}}\left(|q_{2}|+|q_{3}|\right)\leq M^{\prime}(V(q_{2},q_{3}))^{\frac{3}{2}}$
for some $M^{\prime}>0$. Combining this with (3), we obtain
$\begin{array}[]{lll}\dot{V}(q_{2},q_{3})&\geq V(q_{2},q_{3})\left(\kappa-M^{\prime}\sqrt{V(q_{2},q_{3})}\right).\end{array}$
Let $\eta_{*}:=\frac{\kappa^{2}}{4M^{\prime 2}}$, and pick $\eta$ such that $0<\eta\leq\eta_{*}$. Then, for all $p\in\left\{q\in T^{-1}\tilde{\mathcal{B}}_{16}\ |\ V(q_{2},q_{3})<\eta\right\}$, it holds that $\dot{V}(p_{2},p_{3})\geq\frac{\kappa}{2}V(p_{2},p_{3})>0$. Therefore, any solution $q(t)$ with $V(q_{2}(0),q_{3}(0))>\eta$ satisfies $V(q_{2}(t),q_{3}(t))>\eta$ for all $t\geq 0$, and hence ~$\tilde{\mathcal{H}}_{\eta}$ is indeed an invariant set of the dynamics. Note that by the definition of ~$\tilde{\mathcal{H}}_{\eta}$, the equilibrium point 00 is not in ~$\tilde{\mathcal{H}}_{\eta}$ (see also Fig. 1).
@@
Summarizing, there exists $\eta_{*}>0$ such that, for any $0<\eta\leq\eta_{*}$, ~$\tilde{\mathcal{H}}_{\eta}$ is a closed invariant set, and for all initial conditions $q(0)\in\tilde{\mathcal{H}}_{\eta}$ the solution $q(t)$ is bounded and keeps a positive distance from the unique equilibrium 00. Combining this with the strong Poincaré-Bendixson property of the system implies that any trajectory emanating from ~$\tilde{\mathcal{H}}_{\eta}$ converges to a periodic orbit. Since any $q\in T^{-1}\tilde{{\mathcal{B}}}_{16}\setminus\{0\}$ satisfies $q_{2}^{2}+q_{3}^{2}\neq 0$ (see (17)), we conclude that $T^{-1}\tilde{{\mathcal{B}}}_{16}\setminus\{0\}=\bigcup_{0<\eta<\eta_{*}}\tilde{\mathcal{H}}_{\eta}$. This completes the proof of Theorem 1.  ∎
4 Two Case Studies